# Promotion of Scientific Publications on ArXiv and X Is on the Rise and Impacts Citations

### v1

@@ Introduction
Table 1: Confounding variables and the reason to control for the confounding variable. The set of confounders under ArXiv-first effect is common for all treatments. We cumulatively add or remove new confounding variables with a ”+” or ”-” in front of their names, respectively.
Figure 4: When do ArXiv and X or Twitter publications happen? Histogram of the number of days since (a) the conference submission date, (d) the camera-ready deadline to the first ArXiv submission, histogram of the number of days since (b) the conference submission date, (e) the camera-ready deadline to ArXiv revisions, and histogram of the number of days since the (c) the conference submission date, (f) the camera-ready deadline to the first tweet by an author.
Our outcome variables here are the number of total citations (RQ2-4) and the number of highly influential ones (RQ5). We only consider citations for five years since the date of first publication (on ArXiv or in the conference). By doing this, we can compare older papers, that might have higher citation counts as they were published earlier, directly with newer ones, that is we control for the early access. To estimate the treatment effects for our hypotheses, we control for potential confounders. Table 1 details all control variables. First, we describe the core set of control variables, which is common for calculating all the effects (RQ2-4). Then, we describe the adjustment in control variables for calculating the effect of tweeting (RQ4).
@@
The self-selection postulate (Kurtz et al. 2005) states that better papers or papers from established authors are published more on ArXiv. This is referred to as the paper quality bias and the established author bias in this paper, respectively. To control for the established author bias, we use the total citation counts and h-index of the first author and the top author (selected based on maximum total citation counts). We use the top 150 conferences (estimated based on the number of papers in each conference in our dataset) as a one-hot encoded feature. We also control for the average number of citations of papers published per conference, and the conference rating from the GII-GRIN-SCIE Conference Rating Initiative  [^7] https://scie.lcc.uma.es:8443/. These confounding variables give us a signal about the quality of the paper and hence can be used to control for the quality bias.
@@
We also use information about confounding variables like the subject category of the paper other than computer science (like physics, biology etc.) if it exists, the computer science subfield it belongs to (like artificial intelligence, databases, etc.), and the conference it was published in to control for biases arising from certain computer science areas being more popular at different periods of time and hence accruing more citations. We call this area bias. The publication date is also used as a control variable to adjust for temporal changes.
@@
We use standardization and Targeted Maximum Likelihood Estimation (TMLE) using superlearners for estimating average treatment effects. The treatment variable is always binary and is estimated using a superlearner trained on a mixture of a logistic regression model, Discrete Bayesian Additive Regressive Trees (DBARTS) and Generalized Additive Models (GAM). The outcome model to estimate the number of citations is also a superlearner that utilizes Bayesian Additive Regressive Trees (BART), and random forest regression. We use TMLE (Gruber and Van Der Laan 2009) because it is semiparametric and makes minimum assumptions of the distribution of the data . It is also doubly robust and the targeting step in TMLE optimizes the bias-variance tradeoff for the average treatment effect instead of the bias-variance tradeoff for the outcome which is common in other methodologies. We use standardization on the outcome model to estimate the treatment effects for individual papers and understand the heterogeneity of each of the effects across papers.
Figure 5: Percentage of papers that are submitted first and revised on ArXiv, and tweeted on X.
@@ Results
Treatment TMLE Standardization
ATE p-value Cross-val AUC ATE p-value Train Cross-val
$R^{2}$ $R^{2}$ $R^{2}$
All citations
ArXiv-first effect $21.1\pm 17.4$ 0.02 0.14 0.77 $9.0\pm 3.7$ $<10^{-6}$ 0.49 0.19
Revision Effect $18.4\pm 17.6$ 0.04 0.13 0.73 $12.0\pm 3.6$ $<10^{-6}$ 0.48 0.14
Effect of Tweeting $44.4\pm 8.0$ $<10^{-6}$ 0.19 0.97 $11.9\pm 1.9$ $<10^{-6}$ 0.62 0.26
Effect of Author(s) $28.4\pm 11.2$ $<10^{-6}$ 0.18 0.99 $14.8\pm 8.9$ 0.001 0.61 0.24
@@
Highly influential citations
ArXiv-first effect $3.4\pm 2.9$ 0.02 0.07 0.76 $1.4\pm 0.8$ $2\times 10^{-4}$ 0.43 0.10
Revision Effect $2.5\pm 4.0$ 0.21 0.07 0.72 $2.1\pm 0.7$ $<10^{-6}$ 0.44 0.10
Effect of Tweeting $7.6\pm 1.61$ $<10^{-6}$ 0.10 0.98 $2.0\pm 0.4$ $<10^{-6}$ 0.56 0.14
Effect of Author(s) $6.2\pm 2.5$ $<10^{-6}$ 0.11 0.99 $2.6\pm 2.3$ 0.02 0.58 0.12
@@
Table 2: Average treatment effect (ATE) of each of our treatments on all citations and highly influential citations over 5 years after first publication.
Temporal trends in paper promotion (RQ1)
In this section, we discuss the statistical characteristics and temporal trends of papers published in computer science conferences to answer RQ1 based on our rich dataset.
We are interested in understanding how the fractions of papers submitted early or revised on ArXiv change over time. The percentage of papers submitted before the conference to ArXiv has increased from 48.6% in 2013 to 90.1% in 2022 (Fig. 5). In fact, submitting early to ArXiv has become the norm in computer science given the fast-moving nature of the field. The percentage of papers that are tweeted also show a similar trend with an exponential increase in the past few years, reaching 78% in 2022. Conversely, the percentage of papers revised on ArXiv does not exhibit this trend and remains relatively constant with minor fluctuations throughout the years.
We also analyze when researchers make their first submissions to ArXiv and when they submit a revised manuscript on ArXiv. Fig. 4 shows that there are two major peaks corresponding to the periods when researchers make their first submission to ArXiv: one at the conference submission date and the other at the final version due date (camera-ready deadline) (Fig. 4(a), 4(d)). We see a similar trend for revisions, with researchers submitting revised versions of their transcripts much more often around the time of the final version due date than the conference submission deadline (Fig. 4(b), 4(e)), likely after incorporating changes that the reviewers point out in the peer-review process. We also see similar trends for the first tweet by an author(Fig. 4(c), Fig. 4(f)) where there is a peak at the conference submission date and the final version due date but this occurs only because 99.02% of papers tweeted by authors in our dataset receive their initial tweet from an author on the same day as their initial ArXiv submission.
Treatment effects (RQ2-RQ4)
@@
We find statistically significant positive effects for the four treatments using both TMLE and standardization, suggesting that submitting early, revising, tweeting, and the author(s) tweeting lead to a higher number of citations over 5 years after publication. The results are summarized in Table 2.
@@
For the ArXiv-first effect, we notice a positive average treatment effect ($21.1\pm 17.4$ using TMLE and $9.0\pm 3.7$ using standardization), which suggests that submitting to ArXiv before conference deadline on average adds from $9$ to $21$ citations in the first 5 years from the initial publication. It is possible that ArXiv, with its millions of subscribers, provides a visibility exceeding that of conference proceedings. The positive ArXiv-first effect could also indicate a significant first-mover advantage (Newman 2009) where a researcher putting their work out early on ArXiv would get higher number of citations compared to others who wait for the completion of the conference peer-review process, which can take several months, disrupting paper’s timeliness and hotness in a fast-paced research area.
@@
The effect of revisions is also positive ($18.4\pm 17.6$ using TMLE and $12.0\pm 3.6$ using standardization), showing that incorporating feedback from the community makes a paper achieve greater impact. It could also mean that important research papers are revised are improved with time, even post-publication, leading to more citations.
@@
We see a positive treatment effect for tweeting as well ($44.4\pm 8$ using TMLE and $11.9\pm 1.9$ using standardization) which suggests that tweets are an impactful tool for publicizing one’s work in the age of the internet. Research paper showing up on X tend to garner much more citations than papers that are not tweeted.
@@
Among the papers that are tweeted, there is also a positive treatment effect if it is tweeted by an author ($28.4\pm 11.2$ using TMLE and $14.8\pm 8.9$ using standardization), as compared to some other account. This suggests that being active on social media, sharing, and promoting one’s work leads to greater visibility and, ultimately, more citations.
Treatment effects per paper.
Fig. 6 displays the heterogeneous distribution of per-paper treatment effects, calculated through standardization. Here, we could not use doubly robust estimation, since it is designed for average treatment effect estimation. The long-tailed distribution of per-paper treatment effects highlights the diverse nature of these effects, indicating effect sizes vary among papers in our dataset based on individual characteristics. We note that for most papers the effect is positive, although not necessarily significant.
@@
Figure 6: Treatment effects for individual papers using standardization. Black and red corresponsds to all citations and highly influential citaions, respectively. The ATE is marked with a solid line.
@@
In Fig. 7 we plot the effect of each treatment in the first year through the fifth year. We see that the ArXiv-first effect and the revision effect increase over the first three years and then plateau in the fourth and fifth years. More interestingly, the effect of tweeting mirrors this pattern indicating that the effect of tweeting is not short-term (i.e., an increase in the number of citations in the first couple of years and then a gradual decrease) but that it actually leads to long-term impact in the five-year time period.
Treatment effects for influential citations (RQ5)
Highly influential citations have positive but smaller effect sizes, given their lower frequency. In our dataset, the average citations are 70.56, with highly influential citations averaging 10.37, resulting in a ratio of 7:1 approximately. This ratio is mirrored in our results (see 2), indicating that promotion on ArXiv and X equally affects highly influential citations. The treatment effects for highly influential citations exhibit also nearly the same temporal trends as raw citation counts (Fig. 6) and have comparable distributions per paper (Fig. 7).
Figure 7: TMLE Average Treatment Effects (ATE) calculated over the first 5 years after publication for (a) ArXiv-first effect, (b) Revision effect, (c) Effect of tweeting, and (d) Effect of author(s) tweeting. Black and red correspond to all citations and highly influential citations, respectively. Error bars correspond to bootstrapped 95% confidence intervals. Results in Appendix E.
Overall, our hypothesis that this novel measure of scientific impact is less affected by tweeting does not hold.
@@ Discussion
In this paper, we look at current trends in computer science publications and the effects of submitting to ArXiv before the conference, revising on ArXiv, and tweeting about a paper.
We find that the practice of publishing on ArXiv before the conference and tweeting about research has increased exponentially over the past few years (Fig. 5), indicating a significant shift in publication dynamics in computer science. We also find positive effects for ArXiv-first submissions, possibly indicating a first-mover advantage (Newman 2009). This presents a challenge, as authors may publish preliminary results on ArXiv to claim ownership, accumulating more citations than more comprehensive studies on the same topic presented at conferences.
@@
Our study suggests a positive effect of tweeting, i.e., one’s research gets more attention if it is promoted on the X platform highlighting social effects in dissemination of scientific information. This finding suggests that conferences prohibiting ArXiv submissions diminish the paper’s potential impact on the community. Further research should explore whether this policy disproportionately hinders early-career researchers (as suggested by Saxon (2023)), as they are unable to leverage the advantages of promoting their work on platforms like X, where sharing ArXiv preprints is the norm. Saxon (2023) speculates that the accessibility of social media equalizes opportunities for early-career researchers, reducing the influence of institutional prestige.
@@
Although the affordances outlined in this paper enhance the speed of scientific communication and are gaining popularity, they present challenges to both the scientific review process and the assessment of research impact. Our results suggest that ArXiv is becoming a centralized repository of scientific knowledge with millions of monthly users who are on the lookout for new papers on the platform, which may provide the papers on ArXiv more exposure than conference proceedings. On the contrary, unreviewed low-quality pre-prints on ArXiv can potentially spread misinformation. The current peer review process in computer science does not incorporate any form of review for ArXiv publications, even following conference publications, despite the widespread readership of the ArXiv versions of papers. One potential answer to this challenge is to adopt post-publication peer review (Ford 2013; O’Sullivan, Ma, and Doran 2021), where the peer review process takes place after publication, in contrast to the traditional pre-publication peer review system. This approach not only facilitates the rapid dissemination of scientific information but also fosters meaningful discussions among researchers, thereby encouraging collaborations. If the computer science community adopts a system that allows organic and quicker ways for providing feedback, this could lead to new paradigms in scientific publishing, such as the ones enabled by systems such as OpenReview.net, where reviews can be instant, visible publicly, and responded. We anticipate such systems would quicken feedback loops, while at the same time present challenges for science evaluation.
@@
Moreover, we find that citations and more modern measures such as highly influential citations are volatile and can be easily influenced by promotion on ArXiv and X. Therefore, there is a need for measures of research impact that are impervious to these effects. Proposed citation count normalization methods, like field-normalized or network-normalized counts (Waltman 2016; Ke, Gates, and Barabási 2023; Hutchins et al. 2016), offer alternative metrics that are independent of publication year or field. In contrast, altmetrics, such as the ones reported by Almetric.com or PlumX, directly depend on a paper’s social media attention. Our comprehensive analysis reveals the necessity for developing research impact metrics that incorporate the social aspect of sharing scientific information. This concept is also put forth by Ke, Gates, and Barabási (2023), advocating the exploration of network-normalized metrics on a modified citation graph that accounts for social factors in the dissemination of scientific information. Classifying citations as topical (based on topical similarities between the papers) or non-topical (based on social or professional connections between the authors) in the citation graph drawing from research on social ties (Aiello, Schifanella, and State 2014; Grabowicz, Ganguly, and Gummadi 2016; Grabowicz et al. 2013), offers a promising avenue since topical citations are expected to carry more relevance than non-topical ones when analyzing the research impact of a paper. Employing such a socially-informed measure would further enhance the accurate assessment of scholarly influence.
@@
In this paper, we have made a conscientious effort to address numerous confounding variables to minimize biases but there still might exist unaccounted confounding variables which may impact our results. Although we use doubly robust estimators, model specification issues could still introduce bias. The higher mean and median citation counts within WikiCFP, where top-tier conferences are prevalent, might lead to higher effect sizes.

### v2

@@ Introduction
Table 1: Control variables and the reason to control for the variable. The set of control variables used to estimate the ArXiv-first effect is also used when estimating the other effects (let us call this set $S$). We cumulatively add or remove new control variables with a ”+” or ”-” in front of their names, respectively. For instance: when estimating revision effect, we use 1 additional feature (over $S$); when estimating tweet effect, we add 5 additional variables and remove 1 (in comparison to $S$).
This section describes the methods used for treatment effect estimation, utilized control variables and the rationale behind using them, followed by details about feature selection and the models used for treatment effect estimation.
One can estimate a treatment effect by modeling outcome variable, $Y$, and applying standardization, or by modeling treatment variable $A$, and applying inverse probability weighting, while controlling for all confounders $W$ (Hernán and Robins 2020). We refer to the regression model of outcome as $\mathbb{\hat{E}}[Y|a,w]$, where $W$ are control variables. The treatment model corresponds to so-called propensity scores $\hat{P}(A=1|w)$. An estimate of treatment effect relies on the specification of respective model. If model specification is correct, then the estimate is unbiased. Doubly robust methods estimate both the model of treatment and outcome, so we have two shots at using the right model. If one of the models is correct and matches reality, then the estimate is unbiased. To further diminish the probability of model misspecification, for our treatment and outcome models we train a superlearner (Van der Laan, Polley, and Hubbard 2007), a weighted ensemble of multiple machine learning sub-models. Superlearner is trained to provide higher weights to the sub-models that generalize better.
The treatment variables, $A$, are the ones defined in RQ2-4. Our outcome variables here, $Y$, are the number of total citations (RQ2-4) and the number of highly influential ones (RQ5). We consider citations in the first five years since the initial publication (on ArXiv or in a peer-reviewed venue). By doing this, we can compare older papers, that might have higher citation counts as they were published earlier, directly with newer ones, that is we control for the early access.
Finally, we control for multiple confounders, $W$. Table 1 details all control variables. First, we describe the core set of control variables, used for calculating all the effects (RQ2-4). Then, we describe the adjustment in control variables for calculating the effect of tweeting (RQ4).
@@
The self-selection postulate (Kurtz et al. 2005) states that better papers or papers from established authors are published more on ArXiv. This is referred to as the paper quality bias and the established author bias in this paper, respectively. To control for the established author bias, we use the total citation counts and h-index of the first author and the top author (selected based on maximum total citation counts). We use the top 150 conferences or journals (estimated based on the number of papers in each conference in our dataset) as a one-hot encoded feature. We also control for the average number of citations of papers published per conference, and the conference rating from the GII-GRIN-SCIE Conference Rating Initiative [^7] https://scie.lcc.uma.es:8443/. For physics venues, we control for the impact factor of the journal calculated using data from Scimago [^8] https://www.scimagojr.com/. These control variables give us a signal about the quality of the paper and hence can be used to control for the quality bias.
@@
We also use information about control variables like the subject category of the paper other than computer science or physics (like math, biology etc.) if it exists, the computer subfield it belongs to (like artificial intelligence and databases for computer science, and condensed matter and astrophysics for physics), and the venue it was published in to control for biases arising from certain areas being more popular at different periods of time and hence accruing more citations. We call this area bias. The publication date is also used as a control variable to adjust for temporal changes.
@@
Figure 4: Percentage of papers that are submitted first and revised on ArXiv, and tweeted on X for (a) computer science and (b) physics.
Figure 5: When do ArXiv and X or Twitter publications happen for computer science? Histogram of the number of days since (a) the conference submission date, (d) the camera-ready deadline to the first ArXiv submission, histogram of the number of days since (b) the conference submission date, (e) the camera-ready deadline to ArXiv revisions, and histogram of the number of days since the (c) the conference submission date, (f) the camera-ready deadline to the first tweet by an author.
@@
To increase the probability that our estimates of average treatment effects are unbiased, we use doubly-robust Targeted Maximum Likelihood Estimation (TMLE) (Schuler and Rose 2017). TMLE is semi-parametric and makes minimum assumptions about the distribution of the data.
The TMLE package in R allows the usage of superlearners. Superlearners are ensemble models that use $k$-fold cross-validation to weight $n$ base learners. The outputs of the cross-validation sets from $n$ base learners serve as independent variables that are passed on to a metalearner, typically a linear regression model, with true outputs of the cross-validation sets serving as the dependent variable. Following training, the $n$ coefficients of the metalearner determine the weights for each base learner. Our outcome model $\mathbb{\hat{E}}[Y|a,w]$ is a superlearner that utilizes Bayesian Additive Regressive Trees (BART) and random forest regression. TMLE uses an additional treatment model making the approach doubly robust. The treatment variable is always binary and is estimated using a superlearner $\hat{P}(A=1|w)$ trained on a mixture of a logistic regression model, Discrete Bayesian Additive Regressive Trees (DBARTS) and Generalized Additive Models (GAM). After training the treatment and outcome models, TMLE performs an additional targeting step using both the models which optimizes the bias-variance tradeoff for the average treatment effect instead of the bias-variance tradeoff for the outcome which is common in other doubly robust estimators. This leads to a new adjusted expected outcome $\mathbb{\hat{E}}^{*}[Y|a,w]$, which can be used to estimate average treatment effect. All models are initialized using default parameters in R. The superlearner models employ 10-fold cross-validation to find the best weights for each model.
@@ Results
Temporal trends in paper promotion (RQ1)
We are interested in understanding how the fractions of papers submitted early or revised on ArXiv change over time. The percentage of papers submitted before the conference to ArXiv has increased from 48.6% in 2013 to 90.1% in 2022 (Fig. 4(a)) for computer science. In fact, submitting early to ArXiv has become the norm in computer science given the fast-moving nature of the field. The percentage of papers that are tweeted (by authors or any other account) also shows a similar trend with an exponential increase in the past few years for computer science, reaching 78% in 2022. On the contrary, 90.6% of physics papers in 2013 were being submitted to ArXiv before the journal publication and this percentage went up to 93.5% in 2022 (Fig. 4(b)). The percentage of papers tweeted for physics lies in the 80% to 99% range and does not grow fast like for computer science. The percentage of computer science papers revised on ArXiv remains relatively constant with minor fluctuations throughout the years, whereas in physics increases by about 6% between 2013 and 2022.
As we have fine-grained information about dates of computer science conference submissions and camera-ready deadlines, we also analyze when researchers make their first submissions to ArXiv and when they submit a revised manuscript on ArXiv. Fig. 5 shows that there are two major peaks corresponding to the periods when researchers make their first submission to ArXiv: one at the conference submission date and the other at the final version due date (camera-ready deadline) (Fig. 5(a), 5(d)). We see a similar trend for revisions, with researchers submitting revised versions of their transcripts much more often around the time of the final version due date than the conference submission deadline (Fig. 5(b), 5(e)), likely after incorporating changes that the reviewers point out in the peer-review process. We also see similar trends for the first tweet by an author(Fig. 5(c), Fig. 5(f)) where there is a peak at the conference submission date and the final version due date but this occurs only because 99.02% of papers tweeted by authors in our dataset receive their initial tweet from an author on the same day as their initial ArXiv submission.
Outcome Treatment Computer Science Physics
ATE Relative p-value Cross-val AUC ATE Relative p-value Cross-val AUC
ATE $R^{2}$ ATE $R^{2}$
All citations ArXiv-first effect $21.1\pm 17.4$ 0.30 0.02 0.14 0.77 $3.9\pm 1.1$ 0.20 $<10^{-6}$ 0.13 0.75
Revision Effect $18.4\pm 17.6$ 0.26 0.04 0.13 0.73 $4.3\pm 0.9$ 0.22 $<10^{-6}$ 0.13 0.72
Effect of Tweeting $44.4\pm 8.0$ 0.58 $<10^{-6}$ 0.19 0.97 $6.9\pm 3.5$ 0.29 $8\times 10^{-5}$ 0.23 0.99
Effect of Author(s) $28.4\pm 11.2$ 0.32 $<10^{-6}$ 0.18 0.99 $0.8\pm 3.9$ 0.03 0.74 0.24 0.99
@@
Highly influential ArXiv-first effect $3.4\pm 2.9$ 0.36 0.02 0.07 0.76 $0.2\pm 0.1$ 0.21 $2\times 10^{-5}$ 0.14 0.75
citations Revision Effect $2.5\pm 4.0$ 0.27 0.21 0.07 0.72 $0.2\pm 0.1$ 0.21 $<10^{-6}$ 0.15 0.72
Effect of Tweeting $7.6\pm 1.61$ 0.68 $<10^{-6}$ 0.10 0.98 $0.4\pm 0.2$ 0.33 $2\times 10^{-4}$ 0.18 0.99
Effect of Author(s) $6.2\pm 2.5$ 0.48 $<10^{-6}$ 0.11 0.99 $0.3\pm 0.1$ 0.25 $<10^{-6}$ 0.17 0.99
@@
Table 2: Average treatment effect (ATE estimated via TMLE) of each of our treatments on all citations and highly influential citations over 5 years after first publication. “Relative ATE” is the ATE divided by the mean number of citations or highly influential citations for that discipline.
Average treatment effects (RQ2-RQ4)
@@
We find statistically significant positive effects for the four treatments for computer science (CS) using TMLE, suggesting that submitting early, revising, tweeting, and the author(s) tweeting lead to a higher number of citations over 5 years after publication. For physics, we also find positive effects, although effect sizes are smaller. The results are summarized in Table 2.
@@
For the ArXiv-first effect, we notice a positive average treatment effect ($21.1\pm 17.4$ for CS, and $3.9\pm 1.1$ for physics), which suggests that submitting to ArXiv before conference deadline on average adds $21$ citations for CS and $4$ citations for physics in the first 5 years from the initial publication. It is possible that ArXiv, with its millions of subscribers, provides a visibility exceeding that of conference proceedings. The positive ArXiv-first effect could also indicate a significant first-mover advantage (Newman 2009) where a researcher putting their work out early on ArXiv would get higher number of citations compared to others who wait for the completion of the conference peer-review process, which can take several months, disrupting paper’s timeliness and hotness in a fast-paced research area.
@@
The effect of revisions is also positive ($18.4\pm 17.6$ for CS, and $4.3\pm 0.9$ for physics). This result suggests that incorporating feedback from research community leads to greater impact or that important research papers are more likely to be revised with time, even post-publication, leading to more citations.
@@
We see a positive treatment effect for tweeting as well ($44.4\pm 8$ for CS, and $6.9\pm 3.5$ for physics) which suggests that tweets are an impactful tool for publicizing one’s work in the age of the internet. Research paper showing up on X tend to garner much more citations than papers that are not tweeted.
@@
Among the papers that are tweeted for CS, there is also a positive treatment effect if it is tweeted by an author ($28.4\pm 11.2$ for CS), as compared to some other account. This suggests that being active on social media, sharing, and promoting one’s work leads to greater visibility and, ultimately, more citations. This effect is not statistically significant using TMLE for papers in physics.
We observe the same qualitative results for average treatment effects when we use standardization instead of TMLE. One notable exception is that we get a statistically significant positive effect for authors’ tweets for physics ($3.1\pm 1.0$) with this method. The results for standardization can be found in Appendix E.
Figure 6: TMLE Average Treatment Effects (ATE) computed separately for each year over the first 5 years after publication for (circles) standard citations and (crosses) highly influential citations to (a-d) computer science and (e-h) physics papers. Error bars represent 95% confidence intervals. Detailed results in Appendix F.
Treatment effects for influential citations (RQ5)
Highly influential citations have positive but smaller effect sizes, given their lower numbers. The ratio of citations to highly influential citations is approximately 7:1 for computer science and 20:1 for physics. To account for these differences in the number of citations between our two disciplines and types of citations, we compute relative ATE as ATE divided by the number of citations of the respective type for that discipline. Computer science papers average 69.57 citations and 9.37 highly influential citations, while physics papers average 19.20 citations and 0.97 highly influential citations. [^9] The disparities between disciplines appear because WikiCFP provides all required dates only for top computer science venues, whereas our physics dataset contains many more papers from less prestigious venues, such as Physical Review A, B, C, D, and E. Relative treatment effects helps us account for these differences when making comparisons across citation types and disciplines. For instance, the relative ATE is nearly the same for standard and highly influential citations within a discipline (compare relative ATE values between the top and the bottom of Table 2), likely because relatively speaking the effects are the same for standard and influential citations. The treatment effects for highly influential citations exhibit also nearly the same temporal trends as raw citation counts (Fig. 6). Overall, our hypothesis that this novel measure of scientific impact is less affected by tweeting does not hold.
@@
Figure 7: Treatment effects for individual papers using standardization for (a) computer science and (b) physics. The ATE is marked with a solid line.
Differences in effects between computer science and physics (RQ6)
Average treatment effects.
We look at the relative ATE values to compare treatment effects in computer science and physics (compare left and right sides of Table 2). Notably, the differences in relative ATE values are most pronounced for treatments that are affected by social factors such as word-of-mouth spreading: the effect of tweeting (0.58 for computer science vs. 0.29 for physics, a 50% decrease for physics) and the the ArXiv-first effect (0.30 for computer science vs. 0.20 for physics, a 33.3% decrease). On the other hand, the revision effect, which is the least influenced by social factors, shows the smallest decrease for physics (0.26 for computer science vs. 0.22 for physics, only a 15.3% decrease). These results suggest that citations in computer science are more affected by social influence than in physics, in line with the following findings.
Distributions of treatment effects per paper.
To compute per-paper effects, we could not use doubly robust estimation, since it is designed for average treatment effect estimation. The long-tailed distribution of per-paper treatment effects highlights the diverse nature of these effects, indicating relative effect sizes vary among papers in our dataset based on individual characteristics (Fig. 7). Interestingly, computer science treatment effects display a wider spread than for physics. In other words, the effects of treatment on citation numbers is less predictable for computer science publications, possibly due to presence of a stronger social influence mechanism. This result aligns with the seminal experiments of Salganik at el. indicating that social influence in a (digital) market increases unpredictability of product success in that market (Salganik, Dodds, and Watts 2006).
@@
Finally, we study how treatment effects change over years (Figure 6). For computer science, the ArXiv-first and tweeting effects increase in strength over the first four years. Conversely, in physics papers, these effects do not seem to grow, particularly the effect of tweeting. Here, we also notice the same pattern as seen with the relative ATE values, i.e., the most notable variation occurs in treatments influenced by social promotion, i.e., the effect of tweeting (Figure 6(c, g)), the effect of authors tweeting (Figure 6(d, h)), and the ArXiv-first effect (Figure 6(a, e)), compared to the revision effect that is least related to social promotion (Figure 6(b, f)). These findings, along with the prior two results, suggest social influence cascades (Shakarian et al. 2015; Pastor-Satorras et al. 2015) boost citation rates in computer science papers, while being less influential in physics.
@@ Discussion
We find positive effects for ArXiv-first submissions, possibly indicating a first-mover advantage (Newman 2009). This presents a challenge, as authors may publish preliminary results on ArXiv to claim ownership, accumulating more citations than more comprehensive studies on the same topic.
@@
We find a positive effect of tweeting on the number of citations, i.e., one’s research gets more attention if it is promoted on X, highlighting social effects in dissemination of scientific information. This finding suggests that conferences prohibiting ArXiv submissions diminish the paper’s potential impact on the community.
@@
Although the affordances outlined in this paper enhance the speed of scientific communication and are gaining popularity, they present challenges to the scientific review process. Our results suggest that millions of users who are on the lookout for new papers on ArXiv may lead to more exposures than peer-reviewed venues. However, unreviewed low-quality preprints on ArXiv can spread misinformation. One potential answer to this challenge is to adopt post-publication peer review (Ford 2013; O’Sullivan, Ma, and Doran 2021). This approach not only facilitates the rapid dissemination of scientific information but also fosters meaningful discussions among researchers, thereby encouraging collaboration. If the research community adopts a system that allows organic and quicker ways for providing feedback, this could lead to new paradigms in scientific publishing, such as the ones enabled by systems such as OpenReview.net, where reviews can be posted instantly, visible publicly, and responded by authors. We anticipate such systems would quicken feedback loops, while at the same time present challenges for science evaluation.
We also notice significant differences in treatment effects between computer science and physics. These disparities indicate significant differences in the impact of social influence on citations within each field. Notably, citations in computer science exhibit a greater susceptibility to internet-based promotion, such as early submissions to ArXiv and tweeting, compared to physics. Conversely, treatments like revisions on ArXiv, which are not related to word-of-mouth spreading, yield similar effects in both disciplines. Additionally, citations in computer science appear to benefit from social cascading effects, evidenced by increasing treatment effects over yaers, in contrast to physics. The pronounced social influence in computer science would also render the effects of our treatments on citation numbers more unpredictable, as illustrated by the wider distribution of per-paper treatments in Figure 7. Overall, these three findings indicate an increased propensity for social conformity and collective behavior in computer science compared to physics, which may be related to the hotness, hype, and commercialization of research areas such as artificial intelligence. This hypothesis can be explored by future studies.
@@
Citations and their more modern variants, that is highly influential citations, are volatile and can be easily influenced by promotion on ArXiv and X. Therefore, there is a need for measures of research impact that are impervious to these effects. Proposed citation count normalization methods, like field-normalized or network-normalized counts (Waltman 2016; Ke, Gates, and Barabási 2023; Hutchins et al. 2016), offer alternative metrics that are independent of publication year or field. In contrast, altmetrics, such as the ones reported by Almetric.com or PlumX, directly depend on a paper’s social media attention. Our comprehensive analysis reveals the necessity for developing research impact metrics that incorporate the social aspect of sharing scientific information. This concept is also put forth by Ke, Gates, and Barabási (2023), advocating the exploration of network-normalized metrics on a modified citation graph that accounts for social factors in the dissemination of scientific information. Classifying citations as topical (based on topical similarities between the papers) or non-topical (based on social or professional connections between the authors) in the citation graph drawing from research on social ties (Aiello, Schifanella, and State 2014; Grabowicz, Ganguly, and Gummadi 2016; Grabowicz et al. 2013), offers a promising avenue since topical citations are expected to carry more relevance than non-topical ones when analyzing the research impact of a paper. Employing such a socially-informed measure would further enhance the accurate assessment of scholarly influence.
@@
In this paper, we have made a conscientious effort to address numerous control variables to minimize biases but there still might exist unaccounted control variables which may impact our results. Although we use doubly robust estimators, model specification issues could still introduce bias.